OpenStudy (anonymous):

I need help with #5 & #6. My attempts will be in comments.

OpenStudy (anonymous):

!

OpenStudy (danjs):

you have tangent correct so far... the cotangent is 1/tan or adjacent/opposite

OpenStudy (danjs):

the definitions are here on the first page

OpenStudy (anonymous):

would it be 7/5? @DanJS

OpenStudy (danjs):

yes, if it were for that same angle x as before... it says Cot(y) this time

OpenStudy (anonymous):

@DanJS Awesome!! Now, Could you guide me through 6?

OpenStudy (anonymous):

i see whatever cos is i flip it to make it sec

OpenStudy (danjs):

since it is a right triangle, angles x+y=90 notice y=90-x and \[\cot(90-x) = \tan(x)\]

OpenStudy (anonymous):

whoa, okay.

OpenStudy (danjs):

listed under cofunctions on that pdf cot(y) = cot(90-x) = tan(x)

OpenStudy (danjs):

sin and cos involve the hypotenuse, you should figure that length out first

OpenStudy (danjs):

nm, you have it from another prob you did yes, flip cos over, sec(x)=1 / cos(x) however, the secant is for y, not x

OpenStudy (anonymous):

i got 

OpenStudy (danjs):

\[\sec(y) = \frac{ 1 }{ \cos(y) }=\frac{ 1 }{ 5/\sqrt{74} }=\frac{ \sqrt{74} }{ 5 }\]

OpenStudy (danjs):

cos(x) is just adj/hp = 7/root(74)

OpenStudy (anonymous):

Yay! i was about to put does cosx=7/square root 74.

OpenStudy (anonymous):

Now, if you have a second.. did you look over #6 did i do ok on that?

OpenStudy (danjs):

yep, and cos(y) = 5/root(74), so sec(y) is that flipped over

OpenStudy (anonymous):

Sweet!

OpenStudy (danjs):

6 looks blank , but that is what i just did, cos(x) and sec(y)

OpenStudy (anonymous):

omg! lol i meant 4! @DanJS

OpenStudy (danjs):

yes, looks good

OpenStudy (anonymous):

I appreciate it!! You rock!!

OpenStudy (danjs):

welcome, gl